OpenStudy (bj11684):

The total temperature change over 5 days in Pretty City was −15 degrees Celsius. What was the average temperature change each day? −3 degrees Celsius 3 degrees Celsius −10 degrees Celsius 10 degrees Celsius

OpenStudy (rolldextro):

(change)/(time)=-15/5=-3 degrees per day.

OpenStudy (bj11684):

I GAVE U A MEDAL

OpenStudy (rolldextro):

Thanks.